OpenStudy (anonymous):

Differentiate: y= x^(5x)

OpenStudy (anonymous):

lny=5xlnx

OpenStudy (anonymous):

What do I do after?

OpenStudy (gorv):

take log both sides log y = 5x logx differntiate 1/y * dy = 5x*1/x + 5*logx

OpenStudy (anonymous):

okay thanks

OpenStudy (gorv):

1/y * dy = 5x*1/x * dx + 5*logx*dx dy/y=(5+5logx)*dx dy/dx=y*(5+5logx)

OpenStudy (gorv):

but y=x^5x dy/dx=x^5x*(5+5logx)

OpenStudy (anonymous):

okay, yeah I got it, I factored out a 5

OpenStudy (gorv):

okkk :)